OpenStudy (anonymous):

Find the area of the region bounded by the following lines, x^2+y^2=12, xsqrt6=y^2, (x>(y^2/6)) Present a sketch of the region and shade it. Show all working, please

OpenStudy (anonymous):

\[x^2 +y^2 = (2\sqrt{3})^2\] Circle with (0,0) as center and radius \(2\sqrt{3}\)

OpenStudy (anonymous):

\[y^2 = \sqrt 6 x\]ParabolaPlease think of it as Parabola though it doesn't look like one